OpenStudy (anonymous):

Which method is correct when finding the product of (x + 2)2? Justify your answer in complete sentences. plz explain fully!

OpenStudy (anonymous):

You would do (x+2) first. Then you would multiply the sum of (x+2) By 2. [(x +2)2 X =5 (5+2)2 7x2=14\]

OpenStudy (anonymous):

were factoring now solving

OpenStudy (anonymous):

not*